### v2

# Elastic properties of three-dimensional Yukawa or dust-crystals from molecular dynamic simulations

Abstract
I calculate elastic properties of three-dimensional Yukawa or dust crystals using molecular dynamic simulations. The elastic properties are computed by deforming (compressing/expanding) a cubic dust crystal along different directions. The bulk modulus, shear modulus, and Poisson’s ratio are determined as a function of shielding parameter $\kappa$ and strong coupling parameter $\Gamma$. Furthermore, I calculate sound wave velocity and shear wave velocity using bulk and shear modulus values and compare them with the velocities obtained from the slope of longitudinal and transverse wave dispersion relations. It is found that the sound and shear velocities obtained using bulk and shear modulus agree with the experimental values.
I Introduction
Dusty plasmas consist of electrons, ions, neutrals, and immersed dust particles. It is observed in both natural and laboratory settings. In nature, dusty plasmas are observed in Saturn’s ring, interstellar clouds, and cometary tails Shukla and Mamun (2015). In the laboratory, it is present in the plasmas of fusion devices, rocket exhaust, and created in experimental devices under controlled conditions Chu and Lin (1994); Chaubey and Goree (2022). The dusty plasma offers a model system to study generic phenomena such as self-organization and transport at particle level Morfill and Ivlev (2009). Both theoretical and experimental studies have been carried out to study generic phenomena in dusty plasmas, such as crystallization Thomas et al. (1994); Hamaguchi, Farouki, and Dubin (1997); Maity and Das (2019), single-particle dynamics Maity et al. (2018); Deshwal et al. (2022), solitons Kumar, Tiwari, and Das (2017); Bandyopadhyay et al. (2008); Donkó et al. (2020); Kumar and Sharma (2023), shocks Sharma et al. (2016); Arora and Maity (2023), spiral waves Kumar, Patel, and Das (2018); Kumar and Das (2018), vortices Bailung et al. (2020); Dharodi (2020), and Mach cones Samsonov et al. (1999).
A typical micron-sized dust particle carries an electronic charge ranging from approximately -10,000e to -20,000e and possesses a mass approximately $10^{13}$ to $10^{14}$ times that of ions. Dusty plasma can be modeled by a system of point particles interacting through a Yukawa potential, also known as a shielded Coulomb potential, described by the following expression Konopka, Morfill, and Ratke (2000):
${U(r)=\frac{Q^{2}}{4\pi\epsilon_{0}r}\exp(-\frac{r}{\lambda_{D}})}\,,$ (1)
where $Q=-Z_{d}e$ is the charge on a dust particle, $r$ is the separation between two dust particles, and $\lambda_{D}$ is the Debye length of background plasma. The Yukawa system can be characterized in terms of two dimensionless parameters $\Gamma={Q^{2}}/{4\pi\epsilon_{0}ak_{B}T_{d}}$ (known as the strong coupling parameter) and $\kappa={a}/{\lambda_{D}}$ (known as the shielding parameter). Here, $T_{d}$ and $a$ are the dust temperature and the Wigner-Seitz (WS) radius, respectively. The Yukawa inter-particle interaction is also used to model other systems, including charged colloids Palberg et al. (1995); Terao and Nakayama (1999), electrolytes Levin (2002); Lee et al. (2017), and strongly coupled electron-ion plasmas Lyon, Bergeson, and Murillo (2013); Vorberger et al. (2012).
Elastic properties are crucial material characteristics that indicate the stiffness of a material. They are linked to the propagation of waves and their velocities along specific crystallographic directions within a material. Wave propagation and viscoelasticity in dusty plasmas have been extensively investigated through experiments and simulations over the past few decades Morfill and Ivlev (2009). The shear modulus of two-dimensional dusty plasmas has been studied in both experiments and simulations Liu and Goree (2017); Wang, Huang, and Feng (2019). Earlier simulation studies have been carried out to calculate the elastic properties of three-dimensional Yukawa systems Robbins, Kremer, and Grest (1988); Khrapak and Klumov (2020) . However, these properties are not calculated directly by deforming Yukawa or dust-crystals.
In this paper, I calculate the elastic properties of three-dimensional dusty plasma using molecular dynamic simulations by deforming (compressing/expanding) a cubic dust crystal along different directions. The stress and strain of the deformed crystal are used in the calculation of the elastic stiffness constants of dusty plasma. Furthermore, these elastic constants are utilized to calculate the bulk modulus, shear modulus, and Poisson’s ratio. I determine the bulk modulus, shear modulus, and Poisson’s ratio as functions of the shielding parameter $\kappa$ and the strong coupling parameter $\Gamma$. These elastic properties are calculated using experimental dusty plasma parameters. I assess the accuracy of my elastic property calculations by comparing the sound velocity and shear velocity obtained using it with those obtained from the slope of longitudinal and transverse wave dispersion relations.
The paper is organized as follows. Section II provides details of molecular dynamic (MD) simulations. In Section III, I report the results for elastic properties as a function of shielding parameter $\kappa$ and strong coupling parameter $\Gamma$. Section IV contains a brief summary of the study.
II Simulation Method
[…]
The particles are equilibrated at a given temperature using Langevin dynamics, which reads as
$m\ddot{\mathbf{r}_{\mathrm{i}}}=-\sum_{j}\mathbf{\nabla}U_{ij}+\mathbf{F}_{f}+\mathbf{F}_{r}\,,$ (2)
where $U_{ij}$ is interaction potential, $\mathbf{F}_{f}$ is frictional force on the particles, and $\mathbf{F}_{r}$ is random force (kicks) on the particles. I choose the simulation time step of 100 $\mu$s, which ensures a fine discretization along the temporal domain and good resolution of the underlying dust kinetics. The dust particles are evolved for the 40000 time steps for the equilibration of the system. Fluctuations in temperature, pressure, and total energy of the system over time are monitored to verify system equilibration.
Following equilibration, the system is prepared for the calculation of elastic properties under isothermal conditions. The cubic box is transformed into a triclinic box and equilibrated again for 450000 time steps. Subsequently, the triclinic box undergoes deformation along various directions. After each expansion and contraction, the triclinic box is equilibrated for 50000 time steps. The resultant changes in the stress of crystal, namely $P_{x}$, $P_{y}$, $P_{z}$, $P_{xy}$, $P_{xz}$, $P_{yz}$, with deformation (strain) along X, Y, Z, XY, XZ, and YZ directions (Voigt deformation component), respectively, are used in the calculations of elastic stiffness constants.
The convergence of elastic properties with the magnitude of deformation (expansion/compression) is checked carefully. A deformation magnitude of 0.075 (in strain units) is employed in the computations. The average of stresses is taken at 1000 time steps. The finite size effects on elastic properties are examined, which reveals that 10235 atoms are adequate to mitigate such effects.
III Results
The purpose of this study is to calculate the elastic properties of three-dimensional Yukawa or dust-crystal as a function of $\kappa$ and $\Gamma$. Specifically, I compute the bulk modulus, shear modulus, and Poisson’s ratio to characterize the elastic properties. To assess the accuracy of the elastic properties, I calculate sound and shear velocity using them.
III.1 Bulk Modulus
The bulk modulus $K$ is a measure of the resistance of a material to an applied bulk compression. I calculate bulk modulus using elastic stiffness constants via Kittel and McEuen (2018); Robbins, Kremer, and Grest (1988)
$K=\frac{C_{11}+2C_{12}}{3}\,.$ (3)
The bulk modulus at $\Gamma$ = 2000 as a function of screening parameter $\kappa$ is displayed in Fig. 1. It decreases with an increase in the $\kappa$. This reduction occurs because as $\kappa$ increases, the interaction potential length decreases, making compression of the material easier. For large $\kappa$, the Yukawa potential becomes extremely short-range, leading to interactions among particles resembling those in a hard sphere system. The characteristics of bulk modulus at $\kappa$ = 2 with varying strong coupling parameters $\Gamma$ is shown in Fig. 2. The bulk modulus decreases with an increase in $\Gamma$. In this study, $\Gamma$ is increased by reducing the temperature of the dust particles. Consequently, as $\Gamma$ increases, it becomes easier to compress the particles due to reduced thermal pressure. However, at higher $\Gamma$ values, the rigidity driven by the interaction potential overtakes the rigidity driven by temperature, resulting in no further change in bulk modulus with decreasing temperature (see Fig. 2).
Figure 1: Bulk modulus $K$ of dust crystal at the strong coupling parameter $\Gamma$ = $2000$ as a function of shielding parameter $\kappa$.
Figure 2: Bulk modulus $K$ of dust crystal at shielding parameter $\kappa$ = $2$ as a function of strong coupling parameter $\Gamma$.
To validate the accuracy of my bulk modulus calculations, I compute the longitudinal sound velocity $C_{S}$ using it. I utilize the relationship $C_{S}$ = $\sqrt{K/\rho}$ for the calculation of sound velocity Kittel and McEuen (2018). At $\kappa$ = 2.0 and $\Gamma$ = 3000, the sound velocity obtained using bulk modulus is 2.01 cm/s, which closely matches the sound velocity value of 2.15 cm/s obtained from the slope of the longitudinal wave dispersion relation, see Fig. S1 in supplementary material sup . The details of the sound velocity calculation from the dispersion relation are provided in the supplementary material and have also been recently reported by Kumar et al. (2023). The experimentally measured and theoretically calculated values of sound velocity by Bailung et al. (2020) at the same parameters are 4.5 cm/s and 3.06 cm/s, respectively. From these comparisons, it is evident that my bulk modulus values yield correct sound velocity in the dusty plasma medium. The strongly coupled dusty plasmas are extremely soft so that the bulk modulus of solid dusty plasma is much smaller than those of typical solids like metals. For example, the bulk modulus of aluminum at ambient conditions is 79 GPa Simmons (1965).
III.2 Shear Modulus
The shear modulus $G$ is a measure of a material’s resistance to an applied shearing deformation. I calculate the shear modulus using the following formulas Kittel and McEuen (2018); Robbins, Kremer, and Grest (1988):
$G_{1}=C_{44}$ (4)
and
$G_{2}=\frac{C_{11}-C_{12}}{2}\,.$ (5)
Here, $G_{1}$ and $G_{2}$ are the shear modulus along 100 and 110 crystallographic directions, respectively. The shear modulus $G_{1}$ as a function of $\Gamma$ at different $\kappa$ values is displayed in Fig. 3. Additionally, the shear modulus at $\Gamma$ = 2000 as a function of $\kappa$ along the 100 and 110 crystallographic directions are shown in Fig. 4. The values of $G_{1}$ and $G_{2}$ differ at small $\kappa$ values because the particles are in a crystalline phase, resulting in different particle arrangements along the 100 and 110 directions, leading to different shear strains along these directions. However, at high $\kappa$ values, the particles are in a liquid phase and exhibit a uniform distribution along all directions resulting in $G_{1}$ and $G_{2}$ being equal. In the crystalline phase, the shear modulus decreases with an increase in the shielding parameter, which is a direct consequence of the reduction in inter-particle (inter-layer) interactions (see Fig. 3 and Fig. 4). The shear modulus initially decreases with an increase in the strong coupling parameter $\Gamma$, and later it increases with further increases in the strong coupling parameter due to an increase in inter-particle interactions. This increase in inter-particle (inter-layer) interaction results in an increase in resistance to an applied shear deformation. At very high $\Gamma$ values, temperature-driven shear deformation becomes very low and is suppressed by inter-particle driven shear resistance, consequently causing the shear modulus to saturate.
Figure 3: Shear modulus $G_{1}$ of dust crystal at different shielding parameters $\kappa$ as a function of strong coupling parameter $\Gamma$.
Figure 4: Shear modulus of dust crystal along crystallographic directions of 100 and 110 as a function of shielding parameter $\kappa$. The strong coupling parameter $\Gamma$ of the Yukawa crystal is 2000.
At $\kappa$ = 2.0 and $\Gamma$ = 3000, the shear velocity obtained from the shear modulus relation Kittel and McEuen (2018) $C_{T}$ = $\sqrt{G/\rho}$ is 3.61 mm/s and 3.59 mm/s along the 100 and 110 directions, respectively. These values are close to the experimentally measured value of 4.2 mm/s reported by Pramanik et al. (2002). Nunomura, Samsonov, and Goree (2000) have also reported similar values of shear velocity in their experiments. From the slope of the transverse wave dispersion relation, the obtained shear velocity is 5.36 mm/s, see Fig. S2 in the supplementary material sup . From the comparisons, it is clear that my shear modulus values provide the correct shear velocity in the dusty plasma medium. Similar to the bulk modulus, the shear modulus is much smaller than that of typical solids. For instance, the shear modulus for aluminum at ambient conditions is 25.5 GPa Archer et al. (1972).
III.3 Poisson’s Ratio
The Poisson’s ratio $\nu$ quantifies the deformation (compression/expansion) of a material in directions perpendicular to a specific direction of loading. It is defined as the negative ratio of transverse strain to axial strain. I calculate Poisson’s ratio using the following relation
$\nu=\frac{1}{1+\frac{C_{11}}{C_{12}}}\,.$ (6)
In the above expression, $C_{11}$ and $C_{12}$ are the elastic stiffness constants. The Poisson’s ratio at $\Gamma$ = 2000 as a function of $\kappa$ is displayed in Fig. 5. Initially, Poisson’s ratio decreases with an increase in $\kappa$ because transverse strain decreases with the reduction in interaction length of potential. However, above $\kappa$ > 2.5, it increases with an increase in $\kappa$; this occurs because the Yukawa interaction becomes extremely short-range, resulting in interactions among particles resembling those in a hard sphere system, which leads to much less axial strain. The characteristics of Poisson’s ratio at $\kappa$ = 2 as a function of $\Gamma$ is shown in Fig. 6. With an increase in the strong coupling parameter $\Gamma$, Poisson’s ratio decreases, which is a direct consequence of less strain in the transverse direction due to the reduced thermal motion (thermal pressure) of particles at low temperatures.
Figure 5: Poisson’s ratio of dust crystal at strong coupling parameter $\Gamma$ = $2000$ as a function of shielding parameter $\kappa$.
Figure 6: Poisson’s ratio of dust crystal at shielding parameter $\kappa$ = 2 as a function of strong coupling parameter $\Gamma$.
For dusty plasma, the bulk modulus is much larger compared to the shear modulus, so they can be regarded as effectively incompressible, as it is easier to change shape than to compress. Soft materials with similar characteristics are also considered incompressible in nature Jastrzebski (1976). A perfectly incompressible isotropic material has a Poisson’s ratio of exactly 0.5; for dusty plasma, it is approximately 0.5. The incompressible nature of dusty plasma changes with variations in $\kappa$ and $\Gamma$, see Fig. 5 and Fig. 6. For comparison, it is worth mentioning that Poisson’s ratio of natural rubber and aluminum is 0.4999 and 0.33, respectively Mott and Roland (2009).
IV Summary
I have calculated bulk modulus, shear modulus, and Poisson’s ratio of a cubic Yukawa or dusty-crystal using molecular dynamic simulations. These properties have been computed by deforming (compressing/expanding) the dust crystal along different directions. The characteristics of elastic properties as a function of shielding parameter and strong coupling parameter have been presented. The shear modulus of the Yukawa system has also been determined along the crystallographic directions of 100 and 110. Furthermore, I have obtained sound and shear wave velocities using bulk and shear modulus and compared them with the velocities obtained from the slope of longitudinal and transverse wave dispersion relations. It has been found that the sound and shear velocities obtained using bulk and shear modulus agree with the experimental values. It has also been found from my computations that the bulk modulus is much larger than the shear modulus, which effectively renders dusty plasma as incompressible. This is because it is easier to change shape than to compress a dusty plasma system. I think that my findings on elastic properties can also apply to other systems, such as charged colloids, electrolytes, and strongly coupled electron-ion plasmas.
Acknowledgements.
The author thanks Dr. Attila Cangi (CASUS, Görlitz, Germany) and Dr. Svetoslav Nikolov (SNL, Albuquerque, USA) for helpful discussions. The computations were performed on a Bull Cluster at the Center for Information Services and High Performance Computing (ZIH) at Technische Universität Dresden, on the cluster Hemera at Helmholtz-Zentrum Dresden-Rossendorf (HZDR), and on TACC Frontera at University of Texas at Austin.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.